OpenStudy (anonymous):

what is the equation in standard form of the line which passes through (4,-2) and has a slope of -3

OpenStudy (anonymous):

First we note the point A(4,-2) then we use "the point-slope formula"; y-yA=m(x-xA) then we substitute yA=-2;xA=4;m=-3=>y+2=-3(x-4)=>y+2=-3x+12; and the final form is: ->-3x-y+10=0<-